OpenStudy (anonymous):

Abby walked 3/4 mile on Saturday and 7/12 mile on Sunday. How many miles did she walk in all? A. 5/8 mi B. 5/6 mi C. 1 1/3 mi D. 1 1/2 mi

OpenStudy (anonymous):

C

OpenStudy (anonymous):

i was thinking it was d

ganeshie8 (ganeshie8):

you may be correct, how did u get d ? :)

OpenStudy (anonymous):

3/4 convert to 9/12 and add 7/12 = 16/12 reduce to 4/3= 1 1/3

OpenStudy (anonymous):

http://www.mathsisfun.com/fractions_addition.html have fun

OpenStudy (anonymous):

yelp it's D.